OpenStudy (anonymous):

Suppose y varies directly with x and y= 72 when x = 6 . What direct variation equation relates x and y? What is the value of y when x = 10 ?

OpenStudy (xishem):

When two variables vary directly with eachother, it means that the ratio of the two numbers is always the same. \[\frac{x}{y}\]\[\frac{6}{72}=\frac{10}{y}\]Solve for y.